# QuMAB: Query-based Multi-annotator Behavior Pattern Learning

### v1

@@ 7.3. Model Efficiency Analysis
This section evaluates the model efficiency from two aspects: model complexity and inference time. For model complexity, we use the number of parameters as the evaluation metric; for inference time, we compute the average processing time per sample (Zhang et al., 2023c, d, 2022; Zhang, 2024; Sha et al., 2024). To ensure a fair comparison, Table 6 compares the performance of different methods in the image input pipeline (see Figure 2 in the main paper). Specifically, we evaluate the efficiency of different methods on one perspective of the STREET dataset with 10 annotators. It is worth noting that different methods use different backbone networks. To ensure fairness, we standardize the backbone network to ResNet-34 for all methods. In Table 6, we observe that our model has fewer parameters than the other models, while also maintaining competitive average processing time. This validates our claim in the paper that, compared to the baselines that create separate conventional models for each annotator, our architecture demonstrates significant advantages in model efficiency.

### v2

@@ 7.4. Model Efficiency Analysis
This section evaluates the model efficiency from two aspects: model complexity and inference time. For model complexity, we use the number of parameters as the evaluation metric; for inference time, we compute the average processing time per sample (Zhang et al., 2023c, d, 2022; Zhang, 2024a; Sha et al., 2024). To ensure a fair comparison, Table 8 compares the performance of different methods in the image input pipeline (see Figure 2 in the main paper). Specifically, we evaluate the efficiency of different methods on one perspective of the STREET dataset with 10 annotators. It is worth noting that different methods use different backbone networks. To ensure fairness, we standardize the backbone network to ResNet-34 for all methods. In Table 8, we observe that our model has fewer parameters than the other models, while also maintaining competitive average processing time. This validates our claim in the paper that, compared to the baselines that create separate conventional models for each annotator, our architecture demonstrates significant advantages in model efficiency.